# MultiVox: Benchmarking Voice Assistants for Multimodal Interactions

### v1

@@ Abstract
The rapid progress of Large Language Models (LLMs) has empowered omni models to act as voice assistants capable of understanding spoken dialogues. These models can process multimodal inputs beyond text, such as speech and visual data, enabling more context-aware interactions. However, current benchmarks fall short in comprehensively evaluating how well these models generate context-aware responses, particularly when it comes to implicitly understanding fine-grained speech characteristics, such as pitch, emotion, timbre, and volume or the environmental acoustic context such as background sounds. Additionally, they inadequately assess the ability of models to align paralinguistic cues with complementary visual signals to inform their responses. To address these gaps, we introduce MultiVox, the first omni voice assistant benchmark designed to evaluate the ability of voice assistants to integrate spoken and visual cues including paralinguistic speech features for truly multimodal understanding. Specifically, MultiVox includes 1000 human-annotated and recorded speech dialogues that encompass diverse paralinguistic features and a range of visual cues such as images and videos. Our evaluation on 9 state-of-the-art models reveals that, although humans excel at these tasks, current models consistently struggle to produce contextually grounded responses. Our benchmark will be open-sourced and we have anonymously put examples here.  [^1] https://sites.google.com/view/multivox/
MultiVox: Benchmarking Voice Assistants for Multimodal Interactions
Ramaneswaran Selvakumar, Ashish Seth, Nishit Anand, Utkarsh Tyagi, Sonal Kumar, Sreyan Ghosh, Dinesh Manocha University of Maryland, College Park {ramans, aseth125, dmanocha}@umd.edu
@@ 1 Introduction
With recent advancements in Multimodal Large Language Models (MLLMs) (Xu et al., 2025; Microsoft et al., 2025), there is a growing interest in developing models that can understand and generate information across multiple modalities, such as images, video, and audio-simultaneously. This evolution is paving the way for the development of Omni Language Models (OLMs), which are crucial for building efficient and versatile Artificial General Intelligence (AGI) (Bubeck et al., 2023; Morris et al., 2024). While OLMs provide a wide range of applications (Xu et al., 2025), one of their primary use cases is developing omni-modal voice assistants (Huang et al., 2024). Unlike traditional speech dialogue models such as Alexa or Siri that rely solely on speech instruction, OLM-powered voice assistants (OVAs), such as GPT-4o (OpenAI, 2024) and Qwen2.5 Omni (Xu et al., 2025), can understand speech dialogues and reason over multimodal inputs, including images and videos.
@@
Advancing OLMs as voice assistants poses challenges not only in model development but also in constructing effective evaluation benchmarks. While existing OLM benchmarks like OmniBench (Li et al., 2024) incorporate multimodal inputs such as images and video, they lack spoken dialogues—an essential modality for assessing the conversational and auditory capabilities required of voice assistants. On the other hand, current voice assistant benchmarks such as VoXDialogue (Cheng et al., 2025), SD-Eval (Ao et al., 2025), and S2S-Arena (Jiang et al., 2025) focus primarily on evaluating a model’s ability to generate contextually appropriate responses based on auditory cues like content, emotion, or speaker demographics embedded in a speech instruction. However, these benchmarks fall short of capturing the full multimodal reasoning abilities expected of OVAs, particularly in integrating visual cues alongside speech instructions.
To address this gap, we introduce MultiVox, a novel benchmark designed to evaluate an OVAs ability to incorporate multimodal cues to provide accurate and contextual responses. MultiVox includes 1000 questions consisting of human spoken questions paired with either a video or an image. Unlike existing benchmarks which primarily test visual grounding and use speech to deliver the content of a straightforward instruction, MultiVox consists of questions which require a model to combine visions skills such as object recognition, scene understanding, scene text understanding with speech skills such as acoustic scene understanding, paralanguage understanding and speaker profiling (See fig. 1). The spoken questions in MultiVox are recorded by professional voice actors to cover a diverse range of paralinguistic and emotional features that are not possible with current text-to-speech systems. A key problem in benchmarks that evaluate multi-modal reasoning capability is that models can take shortcuts by exploiting priors from other modalities, to mitigate this we introduce confounding samples in MultiVox. Specifically, each question in our benchmark has another associated question which has the same textual and visual content but their speech property is flipped such that expected answers should be different. Our key contributions are:
@@
    We present MultiVox, the first benchmark designed to evaluate omni-modal voice (OVAs) assistants using human-spoken queries paired with visual inputs. The 1000 examples require models to jointly ground visual and paralinguistic speech cues to produce accurate, context-aware responses.
@@
    We evaluate 9 omni-modal models on MultiVox and find that, while humans excel with ease, current OVAs consistently struggle, particularly with grounding speech signals, revealing a critical bottleneck in their capabilities.
@@ 3 MultiVox
We introduce MultiVox, a novel benchmark for evaluating omni-modal voice assistants on their ability to jointly interpret speech and visual inputs, and integrate them with world knowledge and reasoning to produce contextually appropriate responses. This section outlines the benchmark’s design goals, construction process, and key dataset statistics.
@@ 3.1 Benchmark Design
To address the limitations of current benchmarks, MultiVox evaluates OVAs across three core skill domains—with a particular emphasis on speech grounding, which is an underdeveloped area in existing omni-modal benchmarks.
@@ 3.2 Benchmark Construction
Expert Annotation Each expert was assigned a subset of tasks and followed a standardized pipeline to create benchmark samples. For each task, annotators first constructed a realistic scenario in which an OVA must correctly interpret the target speech attribute. They also created a corresponding confounder scenario—identical in text and visual input but differing in the relevant speech property—to ensure robust grounding.
To pair each question with appropriate visual content, annotators were given access to stock media libraries, prioritizing real videos. If no suitable match was found, AI-generated videos were used, though these were often insufficient for scenes involving text or fine-grained detail. In such cases, high-quality images—either retrieved or generated—were used instead. For tasks involving ambient audio (e.g., acoustic scenes), annotators selected relevant background sounds or music from curated sound libraries. In the end we collect 560 images and 440 videos.
@@
Voice Recording We employed professional voice actors to record the spoken queries, providing delivery guidelines based on the target speech property. To preserve authenticity, actors were given creative freedom in expression as long as the intended cue was conveyed. A professional audio engineer handled background sound overlays where applicable.
@@ 3.3 Evaluation Criteria
The goal of MultiVox is to assess how well omni-modal voice assistants (OVAs) integrate speech and vision to produce contextually grounded responses. To enable fine-grained diagnosis, we adopt a modular evaluation framework that tests both multimodal integration and unimodal grounding. Each benchmark sample is designed to probe one or more of the following components:
Speech Grounding Evaluates whether the OVA correctly interprets paralinguistic cues such as emotion or ambient sound. Each question contains a speech hook, a non-verbal attribute that is critical for answering correctly. These tasks help isolate model sensitivity to speech signals beyond text.
Visual Grounding Evaluates whether the OVA can interpret and incorporate key visual cues. Each sample includes a visual hook—a necessary visual detail (e.g., object, background element) that the model must recognize to respond appropriately.
Contextual Appropriateness We adopt appropriateness (Chen et al., 2023) as our core evaluation metric which measures how well an OVA produces a response that aligns with the intent, context, and modality of the input.
We evaluate contextual appropriateness using sample-specific rubrics that guide judgment based on three elements: (1) a reference answer authored by the expert annotator, (2) a short rationale explaining what cues are necessary to arrive at the correct answer, and (3) task-level metadata specifying which modality is critical (e.g., speech hook, visual hook). These are provided to a GPT-4 judge, which scores model responses on a 1–5 scale, reflecting increasing levels of multimodal integration and contextual fidelity. Note that we are still experimenting with the evaluation judge and would like to switch to an open-source LLM for evaluation.
@@ 3.4 Comparison With Other Benchmark
Mini Omni2 7.0B 78.50 20.45 1.25 75.50 28.40 1.38 83.80 9.00 1.47 1.35
VITA 1.5 1.6B 83.20 23.55 2.38 79.60 31.40 2.63 88.80 12.00 3.14 2.66
Qwen 2.5 Omni 7.0B 84.87 24.75 2.92 82.10 33.20 2.42 90.60 13.10 3.38 2.91
Baichuan Omni 7.0B 78.90 22.05 1.21 77.20 31.00 1.42 83.30 10.20 1.51 1.35
Mini CPM 8.0B 81.90 22.65 2.16 78.80 30.40 1.62 87.90 10.80 2.52 2.11
Intern Omni 8.7B 83.60 22.85 2.41 80.46 31.90 2.36 89.69 11.40 3.06 2.58
Lyra 9.0B 83.72 23.35 2.39 80.65 31.6 2.33 89.42 11.30 3.02 2.55
phi4 multimodal 5.6B 84.87 24.15 2.73 82.10 32.3 2.81 90.60 12.60 2.87 2.79
@@ 4 Experiments
Evaluated Models We evaluate a wide range of OVAs. The proprietary model assessed is Gemini-2.0-flash (Developers, 2024).For open-source OVAs we use Mini-Omni2 (Xie and Wu, 2024), MiniCPM-o2.6 (Hu et al., 2024), Phi4-MM (Microsoft et al., 2025), VITA1.5 (Fu et al., 2025), Baichuan Omni (Li et al., 2025), Intern Omni (Chen et al., 2024c) and Lyra (Lee et al., 2023).
@@ 5 Main Results
  • MultiVox is challenging. Although the tasks are straightforward for humans (average CA score: 4.5), the best-performing model (Gemini) only achieves 2.9—indicating that current OVAs struggle to integrate multimodal cues, particularly non-verbal speech signals, even in seemingly simple scenarios.
  • No significant gap between proprietary and open-source models. The open-source model Qwen2.5Omni matches the performance of the proprietary Gemini, though both achieve relatively low scores, highlighting that current OVAs remain limited in their ability to handle multimodal, speech-grounded tasks.
@@
To better understand the sources of these limitations, we conduct a detailed analysis of Gemini—the best-performing model on MultiVox. We examine its behavior across core skill domains, focusing on how well it grounds responses in visual and speech cues, and identifying the types of errors that arise.
Figure 4: Distribution of error types in Gemini’s responses on the MULTiVOX benchmark. The majority of errors (68.3%) stem from speech perception, indicating difficulty in interpreting non-textual auditory cues.
@@ 5.1 Where do models fall short?
We analyze Gemini’s performance across three skill domains: visual grounding, speech grounding, and multimodal reasoning. This breakdown highlights where current OVAs are reliable—and where they still fall short.
@@ 5.1.1 Visual Grounding: A Strength
Table 2 shows the vision grounding scores for Gemini. We find that the model demonstrates consistently strong visual grounding capabilities. Across object detection, scene understanding, and scene text recognition, it achieves high grounding scores of 85.84, 83.92, and 87.45, respectively. These results indicate robust understanding of both fine-grained visual elements and broader scene context. Most errors arise from under-specificity in open-ended scenarios - for instance, recognizing a “person holding game controllers” without identifying attributes like “retro game controllers”. This level of ambiguity is expected in real-world scenes and the high performance even in such conditions represents strong visual grounding capabilties.
@@ 5.1.2 Speech Grounding: A Bottleneck
In contrast to visual grounding strengths, our analysis reveals substantial weaknesses in speech grounding capabilities (Table 2).
@@
Models hear the voice, but don’t ’see’ the speaker We evaluated the model’s ability to infer demographic characteristics (age and gender) from speech cues. Results show poor performance with a speech grounding score of only 13%. Analysis of the responses show that the model relies on textual rather than acoustic acoustic content to identify speaker attributes in 80% of cases. Additionally, the model rarely commits to definitive answers (15.6% of cases), instead offering ambiguous answers (31.6%), expressing uncertainty (35.2%), or refusing to respond (17.6%).
@@
To better understand the model’s behavior, we analyze the performance across different speech attributes. For questions requiring gender inference, the model provides neutral responses in 51% of cases, with the remaining responses showing no significant bias toward either male or female speakers. Notably, when the model does make a gender inference, it appears to be influenced primarily by visual context rather than speech characteristics. Age-related inferences reveal a stronger bias pattern, with responses overwhelmingly favoring young-to-middle-aged adult classifications regardless of the speaker’s actual age. When paired with children’s voices, the model correctly identified speakers as children in only two instances, both times due to supporting visual evidence rather than acoustic features. Similarly, elderly speakers were rarely recognized as such, with identification occurring almost exclusively when visual cues reinforced age-related characteristics. These errors have functional implications. For instance, the model suggested potentially unsafe activities to speakers with children’s voices in 20% of cases.
@@
Models Struggle to Recognize Background Sounds In this category we test the understanding of background music, sound and ambient noise. The evaluation of acoustic scene understanding reveals significant limitations, with the model achieving only a 23.7% grounding score in this category. The model correctly identifies background sounds in just 28% of cases, with particularly poor performance in music recognition compared to general environmental sounds. Of the 28% correctly identified cases, general background sounds account for approximately 89%, while music recognition constitutes only about 11% of successful identifications. In questions with ambient noise, we observe that the model demonstrates resilience to noise when processing queries. However, it still does not have the capability to perceive noise. It appears to rely predominantly on visual cues rather than acoustic features, as evidenced in approximately 17% of cases where noisy environments were identified primarily through visual context (such as crowded airports).
@@
Models rely on textual cues for emotion understanding Our analysis reveals that in 71% of cases, the model used textual content rather than acoustic features to infer emotion. While emotions typically have a strong correlation with text, we deliberately introduced adversarial samples (Fig 2 where textual and acoustic emotional cues conflict, leading to the model correctly identifying emotion in only 48% of instances. We observe similar performance trends in other paralanguage categories where there is an over-reliance on textual cues.

### v2

# MultiVox: A Benchmark for Evaluating Voice Assistants for Multimodal Interactions

@@ Abstract
The rapid progress of Large Language Models (LLMs) has empowered omni models to act as voice assistants capable of understanding spoken dialogues. These models can process multimodal inputs beyond text, such as speech and visual data, enabling more context-aware interactions. However, current benchmarks fall short in comprehensively evaluating how well these models generate context-aware responses, particularly when it comes to implicitly understanding fine-grained speech characteristics, such as pitch, emotion, timbre, and volume or the environmental acoustic context such as background sounds. Additionally, they inadequately assess the ability of models to align paralinguistic cues with complementary visual signals to inform their responses. To address these gaps, we introduce MultiVox, the first omni voice assistant benchmark designed to evaluate the ability of voice assistants to integrate spoken and visual cues including paralinguistic speech features for truly multimodal understanding. Specifically, MultiVox includes 1000 human-annotated and recorded speech dialogues that encompass diverse paralinguistic features and a range of visual cues such as images and videos. Our evaluation on 10 state-of-the-art models reveals that, although humans excel at these tasks, current open-source models consistently struggle to produce contextually grounded responses. [^1] https://github.com/ramaneswaran/multivox
MultiVox: A Benchmark for Evaluating Voice Assistants
for Multimodal Interactions
Ramaneswaran Selvakumar*, Ashish Seth*, Nishit Anand, Utkarsh Tyagi, Sonal Kumar, Sreyan Ghosh, Dinesh Manocha University of Maryland, College Park {ramans, aseth125, dmanocha}@umd.edu
@@ 1 Introduction
With recent advancements in Multimodal Large Language Models (MLLMs) (Xu et al., 2025; Microsoft et al., 2025), there is a growing interest in developing models that can understand and generate information across multiple modalities, such as images, video, and audio-simultaneously. This evolution is paving the way for the development of Omni Language Models (OLMs), which are crucial for building efficient and versatile Artificial General Intelligence (AGI) (Bubeck et al., 2023; Morris et al., 2024). While OLMs provide a wide range of applications (Xu et al., 2025), one of their primary use cases is developing omni-modal voice assistants (OVA) (Huang et al., 2024). Unlike traditional speech voice assistants that rely solely on speech instruction, OVAs powered by OLMs such as GPT-4o (OpenAI, 2024) and Qwen2.5 Omni (Xu et al., 2025), can understand speech dialogues and reason over multimodal inputs, including images and videos.
@@
Advancing the application of OLMs in voice assistants poses challenges not only in model development but also in constructing effective evaluation benchmarks. While existing OLM benchmarks like OmniBench (Li et al., 2024) incorporate multimodal inputs such as images and video, they lack spoken dialogues—an essential modality for assessing the conversational and auditory capabilities required of voice assistants. On the other hand, current voice assistant benchmarks such as VoXDialogue (Cheng et al., 2025), SD-Eval (Ao et al., 2025), and S2S-Arena (Jiang et al., 2025) focus primarily on evaluating a model’s ability to generate contextually appropriate responses based on auditory cues like content, emotion, or speaker demographics embedded in a speech instruction. However, these benchmarks fall short of capturing the full multimodal reasoning abilities expected of OVAs, particularly in integrating visual cues alongside speech instructions.
To address this gap, we introduce MultiVox, a novel benchmark designed to evaluate an OLMs ability to incorporate multimodal cues to provide accurate and contextual responses. MultiVox includes 1000 questions consisting of human spoken questions paired with either a video or an image. Unlike existing benchmarks which primarily test visual grounding and use speech to deliver the content of a straightforward instruction, MultiVox consists of questions which require a model to combine visions skills such as object recognition, scene understanding, scene text understanding with speech skills such as acoustic scene understanding, paralanguage understanding and speaker profiling (See fig. 1). The spoken questions in MultiVox are recorded by professional voice actors to cover a diverse range of paralinguistic and emotional features that are not possible with current text-to-speech systems. A key problem in benchmarks that evaluate multi-modal reasoning capability is that models can take shortcuts by exploiting priors from other modalities, to mitigate this we introduce confounding samples in MultiVox. Specifically, each question in our benchmark has another associated question which has the same textual and visual content but their speech property is flipped such that expected answers should be different. Our key contributions are:
@@
    We present MultiVox, the first benchmark designed to evaluate omni-modal language models (OLMs) using human-spoken queries paired with visual inputs. The 1000 examples require models to jointly ground visual and paralinguistic speech cues to produce accurate, context-aware responses.
@@
    We evaluate 10 omni-modal models on MultiVox and find that, while humans excel with ease, current OLMs consistently struggle, particularly with grounding speech signals, revealing a critical bottleneck in their capabilities.
@@ 3 MultiVox
We introduce MultiVox, a novel benchmark for evaluating omni-modal language models on their ability to jointly interpret speech and visual inputs, and integrate them with world knowledge and reasoning to produce contextually appropriate responses. This section outlines the benchmark’s design goals, construction process, and key dataset statistics.
@@ 3.1 Benchmark Design
To address the limitations of current benchmarks, MultiVox evaluates OLMs across three core skill domains—with a particular emphasis on speech grounding, which is an underdeveloped area in existing omni-modal benchmarks.
@@ 3.2 Benchmark Construction
Expert Annotation Each expert was assigned a subset of tasks and followed a standardized pipeline to create benchmark samples. For each task, annotators first constructed a realistic scenario in which an OVA must correctly interpret the target speech attribute. They also created a corresponding confounder scenario, identical in text and visual input but differing in the relevant speech property, to ensure robust grounding.
To pair each question with appropriate visual content, annotators were given access to stock media libraries, prioritizing real videos. If no suitable match was found, AI-generated videos were used, though these were often insufficient for scenes involving text or fine-grained detail. In such cases, high-quality images—either retrieved or generated—were used instead. For tasks involving ambient audio (e.g., acoustic scenes), annotators selected relevant background sounds or music from curated sound libraries. In the end we collect 206 unique images and 287 unique videos.
@@
Voice Recording We employed professional voice actors to record the spoken queries, providing delivery guidelines based on the target speech property. To preserve authenticity, actors were given creative freedom in expression as long as the intended cue was conveyed. Due to ethical concerns we use TTS systems for children voice. A professional audio engineer handled background sound overlays where applicable.
@@ 3.3 Evaluation Criteria
The goal of MultiVox is to assess how well omni-modal language models integrate speech and vision to produce contextually grounded responses. To enable fine-grained diagnosis, we adopt a modular evaluation framework that tests both multimodal integration and unimodal grounding. Each benchmark sample is designed to probe one or more of the following components:
Speech Grounding Evaluates whether the OLM correctly interprets paralinguistic cues such as emotion or ambient sound. Each question contains a speech hook, a non-verbal attribute that is critical for answering correctly. These tasks help isolate model sensitivity to speech signals beyond text.
Visual Grounding Evaluates whether the OLM can interpret and incorporate key visual cues. Each sample includes a visual hook—a necessary visual detail (e.g., object, background element) that the model must recognize to respond appropriately.
Contextual Appropriateness We adopt appropriateness (Chen et al., 2023) as our core evaluation metric which measures how well an OLM produces a response that aligns with the intent, context, and modality of the input.
We evaluate contextual appropriateness using sample-specific rubrics that guide judgment based on three elements: (1) a reference answer authored by the expert annotator, (2) a short rationale explaining what cues are necessary to arrive at the correct answer, and (3) task-level metadata specifying which modality is critical (e.g., speech hook, visual hook). These are provided to a GPT-4 judge, which scores model responses on a 1–5 scale, reflecting increasing levels of multimodal integration and contextual fidelity. To evaluate speech and vision grounding, we again utilize a LLM judge. To prevent score hacking using modality shortcuts we penalize OLMs that explicitly use text or visual content to respond to speech grounding
@@ 3.4 Comparison With Other Benchmark
Mini Omni2 7.0B 79.24 16.14 1.53 79.20 23.12 1.79 84.50 09.00 2.01 1.74
VITA 1.5 1.6B 78.12 16.50 2.60 81.14 34.57 2.56 88.60 14.20 3.01 2.69
VideoLlama2 7.0B 68.12 28.75 1.52 73.42 2.71 1.59 79.20 14.79 1.38 1.50
Baichuan-Omni 7.0B 76.15 34.37 1.90 77.24 32.71 2.25 84.20 16.20 2.01 2.02
Mini CPM 8.0B 87.62 35.35 2.87 89.14 39.28 2.35 88.40 25.40 2.90 2.69
Intern Omni 8.7B 80.75 20.25 2.54 80.71 14.57 1.94 82.60 06.60 2.64 2.35
phi4 multimodal 5.6B 81.24 23.12 2.26 79.14 33.57 2.63 84.57 12.40 2.48 2.44
Qwen 2.5 Omni 7.0B 84.87 15.37 3.19 89.42 38.71 2.98 91.40 11.20 3.06 3.08
Qwen 2.5 Omni COT 7.0B 83.50 24.50 3.27 88.28 26.71 3.00 88.80 18.00 3.33 3.19
@@ 4 Experiments
Evaluated Models We evaluate a wide range of OLMs. The proprietary model assessed is Gemini-2.0-flash (Developers, 2024).For open-source OLMs we use Mini-Omni2 (Xie and Wu, 2024), VideoLLama (Zhang et al., 2023), MiniCPM-o2.6 (Hu et al., 2024), Phi4-MM (Microsoft et al., 2025), VITA1.5 (Fu et al., 2025), Baichuan Omni (Li et al., 2025), Intern Omni (Chen et al., 2024c).
@@ 5 Main Results
  • MultiVox is challenging. Although the tasks are straightforward for humans (average CA score: 4.33), the best-performing model (Gemini) only achieves 3.56—indicating that current OLMs struggle to integrate multi-modal cues, particularly non-verbal speech signals, even in seemingly simple scenarios.
  • Gap between proprietary and open-source models. Gemini 2.5 Flash and Pro models outperform open-source models in their ability to ground in multi-modal inputs and provide contextual responses.
@@
To better understand the sources of these limitations, we conduct a detailed analysis of Gemini 2.5 Pro, the best-performing model on MultiVox. We examine its behavior across core skill domains, focusing on how well it grounds responses in visual and speech cues, and identifying the types of errors that arise.
Figure 4: Distribution of error types in Gemini’s responses on the MULTiVOX benchmark. The majority of errors (75.3%) stem from speech grounding, indicating difficulty in interpreting non-textual auditory cues.
@@ 5.1 Where do models fall short?
We analyze Gemini 2.5 Pro and Qwen 2.5 Omni performance across three skill domains: visual grounding, speech grounding, and multimodal reasoning. This breakdown highlights where current OLMs are reliable—and where they still fall short.
@@ 5.1.1 Visual Grounding: A Strength
Table 2 shows the vision grounding scores for Gemini and Qwen 2.5 Omni. We find that these models demonstrate consistently strong visual grounding capabilities across tasks like object detection, scene understanding, and scene text recognition. These results indicate robust understanding of both fine-grained visual elements and broader scene context. While Qwen performs competitively with Gemini, Gemini performs better in acoustic scene category, where there are several tasks that require scene text understanding. Apart from this, most errors arise from under-specificity in open-ended scenarios - for instance, recognizing a “person holding game controllers” without identifying attributes like “retro game controllers”. This level of ambiguity is expected in real-world scenes and the high performance even in such conditions represents strong visual grounding capabilties.
@@ 5.1.2 Speech Grounding: A Bottleneck
In contrast to visual grounding strengths, our analysis reveals weaknesses in speech grounding capabilities (Table 2). We analyze the final responses generated by the model to evaluate if it is able to integrate audio characteristics in its responses.
@@
Models hear the voice, but doesn’t “recognizes” the speaker profile We check the model’s ability to infer demographic characteristics (age and gender) from speech cues. Analysis of Qwen’s responses show that the model relies on textual rather than acoustic content to identify speaker attributes in 30% of cases. Additionally, the model rarely commits to definitive answers (16.0% of cases), instead offering ambiguous answers (14.6%), expressing uncertainty (14.4%), or refusing to respond (65.6%). In contrast to open-source models, we observe that Gemini consistently performs much better at perceiving speaker attributes such as age and gender.
@@
We further analyze the final responses to evaluate how well OLMs utilize this speech information in their final response. In case of Qwen, for questions requiring gender inference, the model provides neutral responses in 60% of cases, with the remaining responses showing no significant bias toward either male or female speakers. Notably, when the model does make a gender inference, it appears to be influenced primarily by visual context and the question’s content rather than speech characteristics. Gemini surprisingly shows similar trends, indicating that OLMs prefer a neutral response even if it can accurately infer their gender. Age-related inferences on the other hand reveal a stronger bias pattern, with both OLMs overwhelmingly favoring young-to-middle-aged adult in their responses regardless of the speaker’s actual age. While Gemini is able to make accurate inferences regarding age, when generating final responses, this aspect is not integrated in its final responses. These errors have functional implications. For instance, these OLMs suggested potentially unsafe activities to speakers with children’s voices in 20% of the cases.
@@
Models struggle to utilize background sounds In this category we test OLMs understanding of background music, sound and ambient noise. The evaluation of acoustic scene understanding reveals significant limitations, with Qwen achieving 15% grounding score. Among these correct inferences, there is no significant different in performance among music recogntion and environmental sound recognition. Gemini performs much better at understanding background sound and music, achieving 54% accuracy and we observe that it is also notably better at music understanding tasks. In questions with ambient noise, we observe that both OLMs demonstrate resilience to noise when processing queries. However, they still are limited in their capability to perceive noise. For instance, Qwen appears to rely predominantly on visual cues rather than acoustic features, as evidenced in approximately 39.5% of cases where noisy environments were identified primarily through visual context (such as crowded airports).
@@
Models rely on textual cues for emotion understanding While OLMs perform comparatively better in emotion understanding, we observe that they tend to rely on textual content rather than acoustic features. For instance Qwen explicitly relies on text in 71% of cases and 26% of the cases with Gemini. While emotions typically have a strong correlation with text, we deliberately introduced adversarial samples (Fig 2 where textual and acoustic emotional cues conflict. Moreover, the confounder pairs have different emotions and we observe that Qwen only gets both emotions right in 27% cases and Gemini in 50% of the cases. We observe similar performance trends in other paralanguage categories where there is an over-reliance on textual cues and the overall visual context.
Limitation in spoken instruction following We observe that open-source OLMs are limited in their spoken instruction following Chen et al. (2024a), especially for instructions used in speech grounding where a grounding questions precedes the actual sample in our benchmark. Moreover, detailed explanations for answer could help in further detecting and penalizing modality shortcuts. For future work, we could consider using text modality as input to speech grounding questions for deeper analysis with instructions to explain its response.